OpenStudy (anonymous):

The school supply store buys a gross of pencils for $8.64, which is 6 cents.per pencil.The store sells the pencils for 20 cents each or 3 for 50 cents .If half of the pencils are sold individually and half in a bundle of three, then what is the average profit for pencil?

OpenStudy (anonymous):

answer is 12 dollars 1/3 cents . I don't know how . I could not solve it.can you check one more time.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I' could'nt get your answer !

OpenStudy (anonymous):

the answer is for the average of the profits?

OpenStudy (anonymous):

it's okay.I didn't get it either

OpenStudy (anonymous):

it's from pizitz math tournament 8th grade written test

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

6 cents per pencil

OpenStudy (anonymous):

sorry.it's the buying price

OpenStudy (anonymous):

Did you get it

OpenStudy (anonymous):

72*20 =1440 cents =$14.4 ===> first profit = 14.4 -(72*.06)=$ 10.08 ,,,,,( 72/3)*0.50 =$12 ==> 2nd profit = 12 -(72*0.06)= $ 7.68 ===> average profit = (10.08+7.68)/2 =8.88

OpenStudy (anonymous):

I think my answer here is correct.